The spot price of the market index is $900. A 3-month forward contract on this index is priced at $930. The annual rate of interest on treasuries is 2.4% (0.2% per month). What annualized rate of interest makes the net payoff zero? (Assume monthly compounding.)

A) 4.8%

B) 8.5%

C) 11.2%

D) 13.2%

Answer: D
Answer #1

Spot Price = $900

3-month forward contract on this index is priced at $930.

Let i be the annualized rate of interest that makes net payoff zero

900*(1+i/12)^3 = 930 (Assuming monthly compounding)

or, (1+i/12)^3 = 930/900

or, (1+i/12) = 1.01098

or, i/12 = 0.01098

or i =12*(0.01098) = 0.13176 = 13.176 % = 13.2 %